OpenStudy (anonymous):

f(x)=3/x^2 and g(x)=f^(-1) (x) find the slope of the curve g(x) = f -1(x) at the point (3, 1).

OpenStudy (anonymous):

First, let's find the inverse of \(\frac{3}{x^2}\). We have to solve for \(x\). $$y = \frac{3}{x^2} \\ yx^2 = 3 \\ x^2 = \frac 3 y \\ x = \sqrt{\frac 3 y}$$

OpenStudy (anonymous):

This means \(f^{-1}(x) = \sqrt{\frac 3 y}\). Now, can you take the derivative of that function if I write it like this? $$(3y^{-1})^{1/2}$$

OpenStudy (anonymous):

-1/2(sqrt3)(1/y)^3/2

OpenStudy (anonymous):

$$(3y^{-1})^{1/2}\\ -\frac 1 2 (3y^{-1})^{-1/2} y^{-2}$$ Either you've got it and I'm reading it wrong or you're very close.

OpenStudy (anonymous):

I probably be wrong ha

OpenStudy (anonymous):

Alrighty. Well, what I have is true. So now, we just plug the point x = 3, or y = 3 in this case to find the slope at the point (3, 1).

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Plug y = 3 into \(-\frac 1 2 (3y^{-1})^{-1/2} y^{-2}\) to obtain the answer.

OpenStudy (anonymous):

isn't y=1? at point (3,1)?

OpenStudy (anonymous):

Well, we solved for x. We can treat \(y = x\) now, because \(g(x)\), the inverse, is a function of \(x\).

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

-1/6?

OpenStudy (anonymous):

Yes.